OpenStudy (anonymous):

simplify the expression. (6 - square root of 11)(6 + square root of 11) A. 47 + 12(square root of 11) B. -85 C. 25

OpenStudy (saifoo.khan):

Use FOIL

OpenStudy (kropot72):

Just multiply the bracketted expressions together:\[(6-\sqrt{11})(6+\sqrt{11)}=36-11\]

OpenStudy (kropot72):

The answer is of course 36 - 11 = 25 (Choice C).